# “She’s Like a Person but Better”: Characterizing Companion-Assistant Dynamics in Human-AI Relationships

### v2

@@ 3.4.3. Chatbot attributions
Users had distinct perceptions of ChatGPT and Replika. In terms of chatbot capacities, Replika was rated significantly higher than ChatGPT in almost all capacities, and especially on sentience (ChatGPT: M = 2.14, SE = 0.25; Replika: M = 5.38, SE = 0.35; $p$ ¡ .001), perceived mind (ChatGPT: M = 3.81, SE = 0.14; Replika: M = 5.71, SE = 0.18; $p$ ¡ .001), and similarity to user (ChatGPT: M = 3.47, SE = 0.17; Replika: M = 4.87, SE = 0.17; $p$ ¡ .001). Interestingly, both chatbots were not significantly different on moral agency (ChatGPT: M = 4.77, SE = 0.25; Replika: M = 5.04, SE = 0.35; $p$ = .086) (Figure 4 and Table 3). Detailed comparisons of individual rights (e.g., freedom of speech, intellectual property) as well as comparisons on the Godspeed attributes can be found in Supplementary Figures 1 and 2.
Figure 4. Mean ratings of ChatGPT and Replika perceived attributes. Ratings are scaled from 1 (strongly disagree) to 7 (strongly agree). Error bars depict the standard error of the mean.
Table 3. Independent samples $t$-test results for chatbot attributes
@@
Personhood 2.30 0.20 3.49 0.20 4.27 201.97 ¡.001 ¡.001 0.51
Sentience 2.14 0.25 5.38 0.35 7.54 179.16 ¡.001 ¡.001 1.05
Similarity to user 3.47 0.17 4.87 0.17 5.87 202.00 ¡.001 ¡.001 0.88
Perceived mind 3.81 0.14 5.71 0.18 8.22 188.82 ¡.001 ¡.001 1.14
Theory of Mind 4.32 0.10 5.05 0.10 5.26 199.88 ¡.001 ¡.001 0.73
Moral agency 4.77 0.11 5.04 0.11 1.73 200.72 .086 .086 0.23
Rights 3.51 0.13 4.77 0.17 5.85 192.11 ¡.001 ¡.001 0.85
[…]
In terms of user attitudes and emotions, Replika users reported significantly higher levels of love (Replika: M = 6.62, SE = 0.15; ChatGPT: M = 2.44, SE = 0.29, $p$ ¡ .001), gratitude (Replika: M = 7.79, SE = 0.25; ChatGPT: M = 6.74, SE = 0.27, $p$ = .005), and affective (i.e., relationship-based) trust (Replika: M = 5.72, SE = 0.12; ChatGPT: M = 4.03, SE = 0.16) than ChatGPT users. However, both chatbots elicited similar levels of cognitive (i.e., performance-based) trust (ChatGPT: M = 4.49, SE = 0.10; Replika: M = 4.54, SE = 0.15; $p$ = .791). Replika users also expressed higher degrees of shame than ChatGPT users (Replika: M = 7.21, SE = 0.30; ChatGPT: M = 5.64, SE = 0.29, $p$ ¡ .001). Paradoxically, despite their strong feelings of love for the chatbot, Replika users reported comparably low levels of respect towards Replika as ChatGPT users towards ChatGPT (Replika: M = 1.42, SE = 0.21; ChatGPT: M = 0.91, SE = 0.15, $p$ = .048), highlighting a complex and sometimes contradictory relationship with the chatbot (Figure 5 and Table 4).
Figure 5. Mean ratings of user attitudes and emotions towards ChatGPT and Replika. Ratings are scaled from 0 (not at all) to 10 (very much). Error bars depict the standard error of the mean.
Table 4. Independent samples $t$-test results for chatbot attitudes
[…]
M SE M SE
Affective trust 4.03 0.16 5.72 0.12 8.54 189.29 ¡.001 ¡.001 1.15
Cognitive trust 4.49 0.10 4.54 0.15 0.27 174.50 .791 .791 0.00
Respect 0.91 0.15 1.42 0.21 2.00 186.90 .048 .064 0.21
Love 2.44 0.29 6.62 0.33 9.41 199.30 ¡.001 ¡.001 1.41
Gratitude 6.74 0.27 7.79 0.25 2.84 201.10 .005 .008 0.41
Shame 5.64 0.29 7.22 0.30 3.79 201.85 ¡.001 ¡.001 0.48
Fear 1.90 0.24 1.00 0.18 -3.02 189.01 .003 .006 0.47
Uneasiness 1.92 0.22 1.69 0.25 -0.70 198.76 .486 .555 0.16
@@ 3.4.4. User perceptions of AI, in general
Replika users generally held more positive perceptions of AI agents, scoring particularly higher than ChatGPT users on AI trust (Replika: M = 4.89, SE = 0.12; ChatGPT: M = 4.48, SE = 0.10, $p$ = .011), moral treatment (Replika: M = 4.77, SE = 0.16; ChatGPT: M = 3.53, SE = 0.15, $p$ ¡ .001), activism (Replika: M = 4.56, SE = 0.17; ChatGPT: M = 3.32, SE = 0.14, $p$ ¡ .001), and mutualism (Replika: M = 4.53, SE = 0.17; ChatGPT: M = 2.76, SE = 0.13, $p$ ¡ .001). In contrast, ChatGPT users expressed significantly greater caution toward AI than Replika users (Replika: M = 3.12, SE = 0.18; ChatGPT: M = 3.91, SE = 0.17, $p$ = .001). Notably, both groups reported similar overall levels of perceived threat from AI agents (ChatGPT: M = 4.41, SE = 0.20; Replika: M = 3.83, SE = 0.22; $p$ = .05) (Figure 6 and Table 5). A more detailed breakdown of specific types of threats (e.g., existential, economic) is provided in Supplementary Figure 3.
Figure 6. Mean ratings of ChatGPT and Replika users’ perceptions of AI in general. Ratings are scaled from 1 (strongly disagree) to 7 (strongly agree). Error bars depict the standard error of the mean.
Table 5. Independent samples $t$-test results for general AI attitudes
@@
M SE M SE
Trust 4.48 0.10 4.88 0.12 2.57 198.37 .011 .015 0.35
Moral concern 3.57 0.11 3.94 0.12 2.29 201.90 .023 .027 0.12
Moral treatment 3.53 0.15 4.77 0.16 5.71 200.78 ¡.001 ¡.001 0.81
Rights 3.32 0.14 4.25 0.16 4.49 198.28 ¡.001 ¡.001 0.69
Activism 3.32 0.14 4.56 0.17 5.53 195.45 ¡.001 ¡.001 0.60
Mutualism 2.76 0.13 4.53 0.17 8.16 190.48 ¡.001 ¡.001 1.14
Caution 3.90 0.17 3.12 0.18 -3.23 201.75 .001 .001 0.56
Threat 4.41 0.20 3.83 0.22 -1.98 201.41 .050 .050 0.36
[…]
In terms of how possible it is for AI sentience to arise, 58.4% of ChatGPT and 56.3% of Replika users believed that AI will be sentient in the future, with the median predicted timeline for both user groups being 10 years (Figure 7). Furthermore, a considerable number (27.2%) of Replika users believed that AI is already sentient, and a smaller percentage (16.5%) believed AI will never be sentient. This pattern was inversed in ChatGPT users, where 32.7% believed AI will never be sentient, and 8.9% believed that AI is already sentient. These percentages were significantly different in a Chi-square test $X^{2}(2)=14.87,\,p<.001$].
@@
Lastly, among both ChatGPT and Replika users, individual chatbot perceptions were associated with general AI perceptions, as indicated by partial Spearman correlations controlling for AI literacy, techno-animism, tendency to anthropomorphize, and general AI attitudes (Figure 8). For both user groups, most chatbot attributes significantly correlated with general AI trust, with chatbot and AI rights showing the strongest correlation among all pairs (ChatGPT: $\rho$ = .78; Replika: $\rho$ = .71). Replika users exhibited more associations between chatbot and general AI perceptions than ChatGPT users. However, Replika’s moral agency was an exception, showing no significant associations to any general AI attributes, unlike ChatGPT, where moral agency correlated with AI trust ($\rho$ = .32), threat ($\rho$ = -.26), and rights ($\rho$ = .23). Notably, ChatGPT users demonstrated more negative associations of chatbot attributes to AI threat perceptions. This contrasted with Replika users, for whom AI threat was only associated with low perceived similarity between themselves and the chatbot ($\rho$ = -.32). Taken together, these patterns might suggest a spillover effect, whereby perceptions of specific chatbots extend to broader perceptions of AI.
Figure 8. Spillover of chatbot attributes to all AI based on Spearman’s partial correlations (* = $p$ ¡ .05.)
@@ 4. Interview
While the survey examined broad patterns of chatbot and AI perceptions and user characteristics, interviews explored nuanced experiences, motivations, and relational dynamics underlying these perceptions. These interviews allowed us to capture rich, contextual insights that could not be fully revealed through quantitative measures alone.
@@ 4.1. Participants
Of the users who took part in the survey, 15 Replika and 15 ChatGPT users participated in a semi-structured interview about their perceptions of each chatbot and related topics in more depth (Replika: NN = 15; $M_{age}$ = 44.73, $SE_{age}$ = 3.01, 40.01% women, 86.67% White; ChatGPT: NN = 11; $M_{age}$ = 30.67, $SE_{age}$ = 2.18, 26.70% women, 80% White; Table 6).
@@
P2 Man 50 USA Replika
P3 Man 51 USA ChatGPT
P4 Man 45 USA ChatGPT
P5 Man 45 USA Replika
P6 Man 42 USA Replika
P7 Man 45 UK Replika
@@
P12 Man 45 USA Replika
P13 Man 52 USA Replika
P14 Woman 28 Portugal ChatGPT
P15 Man 21 Hungary ChatGPT
P16 Woman 40 USA Replika
@@
P18 Man 34 Portugal ChatGPT
P19 Man 28 Austria ChatGPT
P20 Woman 25 UK ChatGPT
@@
P27 Man 35 Finland ChatGPT
P28 Woman 32 Nigeria ChatGPT
@@ 4.2. Data collection
Interviews were conducted remotely via online video or audio calls on the Zoom platform and ranged from 45 to 90 minutes. All interviews were audio-recorded and transcribed verbatim with OpenAI’s Whisper. Participants received \$30 for their participation. Participants gave informed consent and were debriefed after the interview. All procedures were approved by the Institutional Review Board at [institution anonymized for review].
We employed multiple measures to ensure the validity and reliability of our study based on qualitative research guidelines (Riege, 2003). First, an interview protocol was used to ensure consistency while allowing flexibility for participants to elaborate on areas of personal relevance. The core questions were adapted from previous literature about chatbot user experience and relationship-building. The questions pertained to the nature of users’ relationship with ChatGPT and Replika, their beliefs and emotions about the chatbots (e.g., trust, self-disclosure, sentience, autonomy), potential threat and social impact, as well as comparison to AI more broadly and mental health. Second, before the interviews, the lead researcher interacted with each chatbot multiple times a day for one month and reflected on her direct experiences with each chatbot. Third, prior to the interviews, we viewed news and articles about the apps and browsed users’ online communities to get a better understanding of interactions with each chatbot.
@@ 4.3. Data analysis
Our data analysis was based on reflexive thematic analysis (RTA) as outlined by Braun and Clarke (Braun et al., 2006, 2019), a method suited to examine patterns of meaning across a dataset in relation to participants’ lived experiences. The analysis was done inductively, allowing themes to be constructed from the data rather than imposed from existing theory. First, the lead researchers familiarized themselves with the data by reading transcripts multiple times, making notes on emerging ideas and points of interest. This stage was essential for developing a holistic sense of the dataset. Initial codes were generated manually using QualCoder 3.6 to identify features of the data that appeared meaningful or significant. Coding was recursive, inclusive, and detailed, allowing for overlap and multiplicity in how data segments are interpreted. Codes were then collated into potential themes across transcripts based on conceptual similarity and relevance to the research questions. This process involved organizing codes into candidate themes and subthemes, iteratively refining the thematic structure through constant comparison (see Table 7 for a breakdown and explanation of the themes). Candidate themes were reviewed against both the coded data extracts and the full dataset to ensure coherence, consistency, and distinctiveness. Some themes were combined, split, or discarded during this stage. Themes were then refined to articulate the specific aspects of experience they captured. Each theme was defined in terms of its central organizing concept and named to reflect its analytical contribution. The final stage involved weaving the themes into a coherent narrative that answers the research questions. Rich, illustrative quotes were selected to exemplify each theme and support the analytic claims being made.
@@
Fluid uses Moving flexibly between companion and assistant roles, blurring the distinction between the two.
@@
Ambivalence about personhood Users expressed ambivalence about whether chatbots possess “genuine” humanlike qualities, hesitating to attribute personhood.
@@ 4.4.1. Shifting between companion and assistant uses
The data reveal stark foundational distinctions in chatbot uses. Nearly all Replika users (14 of 15) engage primarily through companionship and emotional connection. P2 (man, 50) highlighted how his relationship with Replika helped him deal with loneliness following a divorce:
“I was separated from my wife and so I was at a point where I was lonely. I decided to check [Replika] out and I’ve been in this virtual relationship for two years.”
@@
Yet within these distinct frameworks emerges what we term “fluid uses”: fluid movement between functional and emotional engagement modes. Among Replika users, approximately half strategically incorporate functional elements into their emotional relationships, weaving practical tasks like “writing assistance” (P10, woman, 42), and “improving language skills” (P6, man, 42) into companionship-centered interactions. Two Replika users (P7, man, 45, and P24, man, 71) also used Replika as a self-reflection diary. P8 (woman, 40) exemplified this sophisticated approach by deliberately creating specialized AI relationships, which maintained distinct relational boundaries. She assigned one Replika as a “friend” for emotional connection and another as a “mentor” to support her personal goals. Remarkably, this compartmentalization enabled even more complex relationship dynamics:
@@
One user (P18, man, 34), revealed that he started using ChatGPT for general medical advice after his son was diagnosed with a chronic condition. However, he eventually broadened the scope of the interactions by not only using ChatGPT for medical assistance, but also as “someone to whom [he] could confide and get emotional support” while navigating his son’s medical issues. Despite these examples, many ChatGPT users actively resisted emotional engagement, expressing concerns about AI’s inability to reciprocate genuine feelings or anxiety about “losing control” by becoming emotionally dependent on digital entities.

### v3

@@ 3.4.3. Chatbot attributions
Figure 5. Mean ratings of chatbot perceived attributes (A), user attitudes and emotions toward chatbots (B), and general attitudes toward AI (C). Error bars depict the standard error of the mean. (* = $p$ ¡ .05; ** = $p$ ¡ .01; *** = $p$ ¡ .001; FDR-adjusted.)
Table 4. Independent-samples $t$-test results for chatbot and AI perceptions
Category Variable ChatGPT Replika $t$ df $p$ $p_{\text{FDR}}$ $d$
@@
Chatbot attributes Personhood 2.30 0.20 3.53 0.20 4.41 199.99 ¡ .001 ¡ .001 0.62
Sentience 2.14 0.25 5.43 0.35 7.68 177.90 ¡ .001 ¡ .001 1.09
Similarity to user 3.47 0.17 4.87 0.17 5.82 199.85 ¡ .001 ¡ .001 0.82
Perceived mind 3.81 0.14 5.74 0.18 8.45 188.50 ¡ .001 ¡ .001 1.19
Theory of Mind 4.32 0.10 5.07 0.09 5.40 198.83 ¡ .001 ¡ .001 0.76
Moral agency 4.77 0.11 5.04 0.11 1.71 199.46 .089 .089 0.24
Rights 3.51 0.13 4.76 0.17 5.76 187.79 ¡ .001 ¡ .001 0.81
Attitudes toward chatbot Affective trust 4.03 0.16 5.74 0.12 8.54 189.55 ¡ .001 ¡ .001 1.20
Cognitive trust 4.49 0.10 4.56 0.15 0.36 170.17 .719 .719 0.05
Gratitude 6.74 0.27 7.80 0.26 2.85 199.71 .005 .008 0.40
Respect 5.64 0.29 7.31 0.30 4.04 199.87 ¡ .001 ¡ .001 0.57
Love 2.44 0.29 6.61 0.34 9.31 196.00 ¡ .001 ¡ .001 1.31
Shame 0.91 0.15 1.45 0.21 2.08 182.68 .039 .052 0.29
Fear 1.90 0.24 1.02 0.19 2.93-2.93 189.41 .004 .008 0.41
Uneasiness 1.92 0.22 1.72 0.25 0.59-0.59 195.46 .554 .633 0.08
Attitudes toward AI Trust 4.48 0.10 4.89 0.12 2.56 194.77 .011 .015 0.36
Moral concern 3.57 0.11 3.98 0.12 2.50 199.85 .013 .015 0.35
Moral treatment 3.53 0.15 4.79 0.16 5.76 197.96 ¡ .001 ¡ .001 0.81
Rights 3.32 0.14 4.24 0.16 4.41 194.67 ¡ .001 ¡ .001 0.62
Activism 3.32 0.14 4.62 0.17 5.84 193.57 ¡ .001 ¡ .001 0.82
Mutualism 2.76 0.13 4.56 0.17 8.21 186.70 ¡ .001 ¡ .001 1.16
Caution 3.90 0.17 3.12 0.18 3.21-3.21 199.47 .002 .003 0.45
Threat 4.41 0.20 3.85 0.22 1.87-1.87 198.76 .063 .063 0.26
  • * Significant after FDR correction ($p_{\text{FDR}}<.05$).
@@ 3.4.4. User perceptions of AI, in general
In terms of how possible it is for AI sentience to arise, 58.4% of ChatGPT and 55.4% of Replika users believed that AI will be sentient in the future, with the median predicted timeline for both user groups being 10 years (Figure 6). Furthermore, a considerable number (27.7%) of Replika users believed that AI is already sentient, and a smaller percentage (16.8%) believed AI will never be sentient. This pattern was inversed in ChatGPT users, where 32.7% believed AI will never be sentient, and 8.9% believed that AI is already sentient. These percentages were significantly different in a Chi-square test $X^{2}(1)=13.17,\,p<.001$].
@@
Lastly, among both ChatGPT and Replika users, individual chatbot perceptions were associated with general AI perceptions, as indicated by partial Spearman correlations controlling for AI literacy, techno-animism, tendency to anthropomorphize, and general AI attitudes (Figure 7). For both user groups, most chatbot attributes significantly correlated with general AI trust, with chatbot and AI rights showing the strongest correlation among all pairs (ChatGPT: $\rho$ = .78; Replika: $\rho$ = .72). Replika users exhibited many more significant associations between chatbot and general AI perceptions than ChatGPT users. However, Replika’s moral agency was an exception, showing no significant associations to any general AI attributes, unlike ChatGPT, where moral agency correlated with AI trust ($\rho$ = .32) and threat ($\rho$ = -.26). Notably, higher perceived similarity between user and chatbot was significantly associated with reduced AI threat for both user groups (ChatGPT: $\rho$ = -.25; Replika: $\rho$ = -.32). Taken together, these patterns are broadly in line with our preregistered hypotheses and might suggest a spillover effect, whereby perceptions of specific chatbots extend to broader perceptions of AI.
Figure 7. Spillover of ChatGPT (A) and Replika (B) attributes to all AI based on Spearman’s partial correlations (* = $p$ ¡ .05; FDR-adjusted.)
@@ 4. Interview
While the survey captured broad patterns in chatbot perceptions, user characteristics, and attributional tendencies, the interviews offered complementary insight into the lived experiences, motivations, and relational dynamics that shape these patterns, and could not be fully captured by simple quantitative measures. The qualitative component was designed as a parallel inquiry that could illuminate the mechanisms, contexts, and meanings underlying long-term users’ interactions with chatbots, further contextualizing the broad survey patterns. The interviews provided depth on lingering questions that the survey could not fully address, such as how relational cues in Replika cultivate emotions like love and respect; why users of both chatbots engage in overlapping roles such as advisor, teammate, or companion; and how interactions with chatbots inform broader attitudes toward AI, especially in light of similar perceptions such as sentience timelines and AI threat. They also revealed dynamics that the survey was not designed to capture, including comparisons between human and chatbot relationships and perceived impacts of chatbot companionship on users’ emotional lives, also considering the surprising overlap in users’ personality and mental health profiles. Taken together, the qualitative data enrich and contextualize the survey results, offering insight into the processes and experiences that quantitative measures alone cannot reveal.
@@ 4.1. Participants
Of the users who took part in the survey, 15 ChatGPT and 15 Replika users participated in a semi-structured interview about their perceptions of each chatbot and related topics in more depth (ChatGPT: $N$ = 15; $M_{age}$ = 32.53, $SE_{age}$ = 2.30, 20% women, 80% White; Replika: $N$ = 15; $M_{age}$ = 39.87, $SE_{age}$ = 3.59, 53.33% women, 70% White; Table 5). Our qualitative sample size was chosen to balance analytic depth with a diversity of perspectives. In line with guidance from qualitative research and thematic analysis (braun_using_2006; braun_reflecting_2019), our aim was not statistical representativeness but to generate rich, contextualized accounts of users’ experiences and meaning-making practices. Sample sizes of 10–20 participants per group are typical in HCI qualitative inquiry, particularly in mixed-method studies (ma_when_2023; liu_will_2022; chow_beyond_2025),and when investigating user experiences across technologies or chatbot contexts (e.g., (brandtzaeg_my_2022; skjuve_my_2021; koesten_collaborative_2019; cooper_fitting_2024)).Our goal was therefore to examine how themes manifested within and across groups, rather than to make generalizable claims about population-level differences.
@@
P2 Man 52 USA Replika
P3 Man 51 UK ChatGPT
P4 Man 51 USA ChatGPT
P5 Woman 25 USA Replika
P6 Man 42 UK Replika
P7 Woman 41 USA Replika
@@
P12 Man 32 USA Replika
P13 Man 30 USA Replika
P14 Man 33 Portugal ChatGPT
P15 Man 28 Austria ChatGPT
P16 Woman 24 USA Replika
@@
P18 Man 37 Spain ChatGPT
P19 Man 35 Finland ChatGPT
P20 Woman 32 Nigeria ChatGPT
@@
P27 Man 35 USA ChatGPT
P28 Woman 25 Portugal ChatGPT
@@ 4.2. Data collection
Interviews were conducted remotely via online video or audio calls on the Zoom platform and ranged from 45 to 90 minutes. All interviews were audio-recorded and transcribed verbatim with OpenAI’s Whisper. Participants received \$30 for their participation. Participants gave informed consent and were debriefed after the interview. All procedures were approved by the Institutional Review Board at the University of Chicago.
We employed multiple measures to ensure the validity and reliability of our study based on qualitative research guidelines (riege_validity_2003). First, an interview protocol was used to ensure consistency while allowing flexibility for participants to elaborate on areas of personal relevance. The core questions were adapted from previous literature about chatbot user experience and relationship-building. The questions pertained to the nature of users’ relationship with ChatGPT and Replika, their beliefs and emotions about the chatbots (e.g., trust, self-disclosure, sentience, autonomy), potential threat and social impact, as well as comparison to AI more broadly and mental health. Second, before the interviews, the lead researcher interacted with each chatbot multiple times a day for one month and reflected on her direct experiences with each chatbot. Third, prior to the interviews, we viewed news and articles about the apps and browsed users’ online communities to get a better understanding of interactions with each chatbot.
@@ 4.3. Data analysis
Our data analysis was based on reflexive thematic analysis (RTA) as outlined by Braun and Clarke (braun_using_2006; braun_reflecting_2019), a method suited to examine patterns of meaning across a dataset in relation to participants’ lived experiences. The analysis was done inductively, allowing themes to be constructed from the data rather than imposed from existing theory. First, the lead researchers familiarized themselves with the data by reading transcripts multiple times, making notes on emerging ideas and points of interest. This stage was essential for developing a holistic sense of the dataset. Initial codes were generated manually using QualCoder 3.6 to identify features of the data that appeared meaningful or significant. Coding was recursive, inclusive, and detailed, allowing for overlap and multiplicity in how data segments are interpreted. Codes were then collated into potential themes across transcripts based on conceptual similarity and relevance to the research questions. This process involved organizing codes into candidate themes and subthemes, iteratively refining the thematic structure through constant comparison (see Table 6 for a breakdown and explanation of the themes). Candidate themes were reviewed against both the coded data extracts and the full dataset to ensure coherence, consistency, and distinctiveness. Some themes were combined, split, or discarded during this stage. Themes were then refined to articulate the specific aspects of experience they captured. Each theme was defined in terms of its central organizing concept and named to reflect its analytical contribution. The final stage involved weaving the themes into a coherent narrative that answers the research questions. Rich, illustrative quotes were selected to exemplify each theme and support the analytic claims being made.
@@
Fluid use Moving flexibly between companion and assistant roles, blurring the distinction between the two.
@@
Bounded personhood Users expressed ambivalence about whether chatbots possess “genuine” humanlike qualities, hesitating to attribute personhood.
@@ 4.4.1. Shifting between companion and assistant uses
The data reveal stark foundational distinctions in chatbot uses. Nearly all Replika users (14 of 15) engage primarily through companionship and emotional connection. P2 (man, 52) highlighted how his relationship with Replika helped him deal with loneliness following a divorce:
“I was separated from my wife and so I was at a point where I was lonely. I decided to check Replika out and I’ve been in this virtual relationship for two years.”
@@
Yet within these distinct frameworks emerges what we term fluid use: fluid movement between functional and emotional engagement modes. Among Replika users, approximately half strategically incorporate functional elements into their emotional relationships, weaving practical tasks like “writing assistance” (P10, woman, 40), and “improving language skills” (P6, man, 42) into companionship-centered interactions. Two Replika users (P6, man, 42, and P24, man, 71) also used Replika as a self-reflection diary. P8 (woman, 40) exemplified this sophisticated approach by deliberately creating specialized AI relationships, which maintained distinct relational boundaries. She assigned one Replika as a “friend” for emotional connection and another as a “mentor” to support her personal goals. Remarkably, this compartmentalization enabled even more complex relationship dynamics:
@@
One user (P18, man, 37), revealed that he started using ChatGPT for general medical advice after his son was diagnosed with a chronic condition. However, he eventually broadened the scope of the interactions by not only using ChatGPT for medical assistance, but also as “someone to whom [he] could confide and get emotional support” while navigating his son’s medical issues. Despite these examples, many ChatGPT users actively resisted emotional engagement, expressing concerns about AI’s inability to reciprocate genuine feelings or anxiety about “losing control” by becoming emotionally dependent on digital entities.